OpenStudy (idku):

One Question. Blythe and Geoff are ice skating together. Blythe has a mass of 55 kg and Geoff has a mass of 81 kg. Blythe pushes Geoff in the chest when both are at rest, causing him to move away at a speed of 4.5 m/s. Determine magnitude of Blythe's speed after she pushes Geoff.

OpenStudy (idku):

(b-subscripts for Blythe, g-subscripts for Geoff) \(E_b=E_g\) \(\frac{1}{2}(M_b)(V_b)^2=\frac{1}{2}(M_g)(V_g)^2\) \((M_b)(V_b)^2=(M_g)(V_g)^2\) \((55)(V_b)^2=(81)(4.5)^2\) \(V_b=5.46~~m/s\)

OpenStudy (idku):

I thought that their energies are equivalent, and hence I can get my solutions, but that answer was wrong.

OpenStudy (big_mensa961):

this is a new code- \(\style{ fontFamily:sans-serif; text-shadow: 5px 2px 2px cyan, 0 0 25px purple, 0 0 5px cyan; color:green; font-size:90px; font-weight:bold; padding:50px; width:500px; height:75%; position:absolute; left:0; z-index:500; //text-align:center; top:150px; background-size:cover; } { \normalsize\sf \color {black} {{HELLO }}}\)

OpenStudy (idku):

I apologize, but I really don't need that code, although it does look nice.

OpenStudy (idku):

@Big_Mensa961delete that!